Two friends decide to expand how many video games they own. Anna currently owns ten games and every month she will buy another four games to add to her collection. Joan currently owns three games and every month she is going to double her current number of games. After how many months will Joan have a larger collection than Anna


1
Expert's answer
2021-02-16T14:57:46-0500

Anna currently owns games = 10

Anna add games to her collection = 4 games per month

Joan currently owns games = 3

Joan add games to her collection = double the current number of games he owns per month

so collection of Anna : -------------

after 1st month : 10+4 = 14

after 2nd month : 14+4 = 18

after 3rd month : 18+4 = 22

after 4th month : 22+4 = 26


and collection of Joan : ----------

after 1st month : 3*2 = 6

after 2nd month : 6*2 = 12

after 3rd month : 12*2 = 24

after 4th month : 24*2 = 48


so from the above we can see that Joan owns the more number of games than Anna in 3 months.